OpenStudy (anonymous):

Kaera scored 80% on a spelling test with 25 words. How many words did Kaera spell correctly?

OpenStudy (anonymous):

please help, there will be medals.

OpenStudy (mertsj):

To find the number of words spelled correctly, multiply .80 by 25

OpenStudy (anonymous):

20

OpenStudy (mertsj):

Good job!! 20 right out of 25 total = 20/25 = 4/5 = 80%

OpenStudy (anonymous):

It is that easy? thanks!